OpenStudy (anonymous):

How do you draw y=8-3cos4x? Please help!

OpenStudy (goformit100):

@ParthKohli

OpenStudy (helder_edwin):

do you know the graph of cos x?

OpenStudy (anonymous):

It's helpful if we know the following equation: \[y=a \cos (k(x-p)) + v\] Where a = amplitude; p = phase shift; v= vertical shift; and we use k to find the period... period= 2pi/k So here, v=8; a=3; k=4; p=0; period= pi/2 (Remember we use k to find the new period = 2pi/4 = pi/2) The NEGATIVE SIGN in front of the cos flips the graph about the x-axis.

OpenStudy (anonymous):

This is a graph of y=cosx, notice how the amp. here is 1 and the period is 2pi

OpenStudy (anonymous):

This is graph of y=3cosx, notice how the a=3 changes the graph's height from the previous graph (y=cosx)

OpenStudy (anonymous):

This graph (y=8+3cos4x) shows how the vertical shift v=8 changes the graph by moving the entire graph 8 units up *****I made a mistake with the post above this one, the graph it shows is y=3cos4x

OpenStudy (anonymous):

FInally this graph shows how the function does a vertical reflection because of the negative sign in front of the 3